OpenStudy (anonymous):

x2-4/2-x whats the simplified rational expression?

OpenStudy (kropot72):

x^2 - 4 = (x+2)(x-2) ((x+2)(x-2)/(2-x) = -((x + 2)(x - 2))/-(2 - x) = -((x + 2)(x - 2))/(x - 2) = 2 - x

OpenStudy (kropot72):

Sorry, I made a mistake. All the working is correct except the answer is -(x + 2) or alternatively -x - 2.